Calculate pH when: solution prepared by mixing 112.0mL of 1.3x10^-3 M HCL and 444.0 mL of M HClO4 (assume that volumes are additive)

Solution prepared by mixing 1120 mL of 1.3 10-3 M HCl and 444.0 mL of 1.0-103 MHCIO, (Assume that volumes are additive ) Express your answer using two decimal places. pH- Submt Request Answer ide Feedback
Answer #1

112.0 ml of 1.3 * 10^-3 M HCl = 0.112 * 1.3 * 10^-3 = 1.456 * 10^-4 mole.

444 ml of 1.0 * 10^-3 M HClO4 = 0.444 * 1.0 * 10^-3 = 4.44 * 10^-4 mole.

total mole of H3O+ = (1.456 * 10^-4 + 4.44 * 10^-4) = 5.896 * 10^-4 mole.

total volume of the mixture = (444 + 112) = 556 ml.

both HCl and HClO4 are strong acids.

[H3O+] = 5.896 * 10^-4 * 1000 / 556 = 1.06 * 10^-3 M

PH = - log (1.06 * 10^-3) = 2.97

PH = 2.97